# Low-Rank Adaptation for Foundation Models: A Comprehensive Review

### v1

@@ 4.5 LoRA for Long Sequence Modeling
The ability to process long sequences is crucial for many tasks across various domains handled by foundation models [125, 126, 127]. However, standard foundation models are typically constrained by their maximum context length due to the quadratic computational complexity of self-attention with respect to sequence length. To address this limitation, several LoRA-based have been proposed to extend the context window of foundation models.
@@ 4.6 LoRA Serving Systems
S-LoRA [129] introduced a unified paging mechanism to manage both KV cache and LoRA weights in GPU memory, enabling concurrent serving of thousands of LoRA adapters. Punica [130] developed a custom CUDA kernel, Segmented Gather Matrix-Vector Multiplication (SGMV), facilitating efficient batching of requests across different LoRA models on a single GPU. CARASERVE [131] adopted a CPU-assisted strategy, initiating prefill computations for newly requested adapters during GPU loading to mitigate cold-start latency. Furthermore, CARASERVE introduced a rank-aware scheduling algorithm for optimized request routing in multi-GPU clusters. CA-LoRA [132] incorporated LoRA knowledge inheritance and model knowledge recovery mechanisms to maintain performance on personal devices. These innovations enhance the scalability and efficiency of multiple LoRA serving across different computing environments.
5 Application
The effectiveness and efficiency of LoRA in fine-tuning foundation models has led to its widespread adoption across various domains, including language processing, computer vision, speech recognition, multi-modal, code engineering, scientific discovery, recommender systems, graph learning, spatial-temporal forecasting, etc.
@@ 5.1 LoRA in Language Tasks
Language foundation models, such as LLaMA [4], RoBERTa [21] and DeBERTa [133], serve as important base models in LoRA study and have been extensively studied in different language tasks, including natural language understanding [47, 45, 48], question answering [36, 134], machine translation [49, 45], reasoning [56, 49, 75, 135] and natural language generation [55, 61, 62], which can also be seen from Table. This section explores its applications in specialized NLP domains.
Multilingual Language and Dialects Processing. LoRA enables efficient multilingual adaptation while preserving base models’ capabilities through minimal parameter updates. FedLFC [123] introduces a communication-efficient federated learning framework that uses LoRA for multilingual modeling while maintaining the base model’s weights and only updating lightweight LoRA parameters across language families. Similarly, FedHLT [122] employs LoRA with a hierarchical language tree strategy to enable efficient adaptation across dialects while minimizing communication overhead in federated learning settings. LAMPAT [136] leverages LoRA for unsupervised multilingual paraphrasing by applying adversarial training to generate diverse outputs while preserving semantic meaning across languages. HyperLoRA [137] develops a novel approach that generates dialect-specific LoRA adapters using linguistic feature vectors, enabling zero-shot adaptation to unseen English dialects without requiring dialectal training data.
Medical and Clinical Text Processing. Medical and clinical text processing faces unique challenges due to limited data availability and sensitive nature, especially in clinical datasets. LoRA has emerged as a promising solution to address these limitations. Ji et al. [138] utilized LoRA to enhance assertion detection in clinical settings, effectively improving classification accuracy with minimal data. Le et al. [139] also demonstrated that LoRA could effectively adapt pre-trained models for clinical NLP tasks, particularly in data-limited environments. Liu et al. [92] introduced a multi-task framework combining LoRA with MOE to address data imbalance in clinical multi-task scenarios. Shi et al. [140] utilized LoRA in their MedAdapter framework for effective test-time adaptation, thereby avoiding extensive computational resources or sharing data with third parties. Christophe et al. [141] conducted a comparative study with their proposed Med42 model, where LoRA fine-tuning surpassed traditional methods on key medical benchmarks like the USMLE.
@@ 5.2 LoRA in Computer Vision
LoRA has been effectively applied to various vision foundation models, like ViTs [145], DinoV2 [146], MAE [6], SAM [7], and Florence [147] enhancing their adaptability across multiple visual tasks including visual understanding, and visual generation.
@@ 5.2.1 Visual Understanding
Visual understanding contains a broad spectrum of tasks including domain adaptation, semantic segmentation, and content authenticity checking.
@@
ExPLoRA [148] extends pre-trained ViTs by incorporating LoRA modules within self-attention mechanisms, effectively modeling domain-specific style variations, such as those found in satellite imagery, in a low-dimensional space. Similarly, MeLo [149] demonstrates the efficacy of LoRA-based fine-tuning for medical imaging applications, including thoracic disease classification tasks. In addition, Kong et al. [150] apply LoRA to enhance the generalization of vision transformers for face forgery detection across different manipulation techniques and datasets.
A notable advancement in this direction is ConvLoRA [151], which extends the LoRA paradigm to convolutional neural networks for unsupervised domain adaptation in medical image. The architecture combines trainable low-rank decomposition matrices with adaptive batch normalization, establishing a more robust framework for domain transfer.
Semantic Segmentation. The adaptation of visual foundation models, particularly the Segment Anything Model (SAM) [7], has witnessed significant progress in semantic segmentation through LoRA integration. ConvLoRA [151] enhances SAM [7] for semantic segmentation in remote sensing, medical and agricultural images. Building upon this foundation, SAMed [152] demonstrates the effectivness of LoRA-based fine-tuning for multi-organ segmentation tasks. SurgicalSAM [153] applies similar techniques to the domain of robotic surgical instrument segmentation.
Content Authenticity Checking. Detecting synthesized content has become increasingly important with the advancement of generative models. CLIPMoLE [154] presents an architecture that adapts ViT blocks through a hybrid approach, combining shared and separate LoRA modules within an MoE framework. This design enables efficient fine-tuning of CLIP ViT models for transferable image detection across diverse generation techniques. Similarly, MoE-FFD [155] integrates LoRA with convolutional adapter modules while maintaining a frozen ViT backbone, specifically targeting face forgery detection.
@@ 5.2.2 Visual Generation
Image Stylization. LoRA has emerged as a key technique for image stylization tasks due to its ability to rapidly adapt diffusion models to specific artistic styles while preserving the diverse generation capabilities of base model. [159] utilize LoRA to efficiently fine-tune Stable Diffusion [10] for comic-style transfer, specifically adapting it to the style of Calvin and Hobbes comics. Frenkel et al. [160] introduce B-LoRA, which leverages the architecture of Stable Diffusion XL [161] to implicitly separate style and content components of a single image. B-LoRA achieves style-content separation that enables various image stylization tasks, including image style transfer and text-based image stylization. Borse et al. [162] propose FouRA, which performs low-rank adaptation in the frequency domain rather than directly on model weights. FouRA addresses issues of distribution collapse and data copying that can occur when fine-tuning diffusion models with standard LoRA.
Multi-Concept Customization. Another common motivation for using LoRA in visual generation is to address the challenge of generating complex images involving multiple subjects or styles. Shah et al. [163] propose ZipLoRA, employing embedding-decomposed LoRA and gradient fusion to merge independently trained LoRAs without weight manipulation, preserving concept identities. Gu et al.[164] introduce Mix-of-Show, which utilizes LoRA switching and composition during the diffusion process. Similarly, Zhong et al.[165] investigate LoRA switching and composite approaches, alternating between or integrating LoRAs during de-noising to improve multi-element generation. Yang et al.[166] present LoRA-Composer, a training-free method integrating multiple LoRAs with concept injection and isolation constraints to mitigate issues like concept vanishing and confusion. Po et al. [167] propose orthogonal adaptation, which encourages customized weights of different concepts to be orthogonal, enabling efficient merging of independently fine-tuned models without compromising fidelity.
@@
Whether addressing multi-concept synthesis, high-resolution tasks, style transfer, or continual learning, LoRA provides a lightweight solution that enables models like Stable Diffusion and CLIP to maintain high performance while being adapted for new and complex tasks.
@@ 5.3 LoRA in Speech Recognition
LoRA has seen significant use in speech recognition tasks, particularly in efficiently fine-tuning foundation models like Wav2vec2 [8] and Whisper [170].
In fake audio detection, Wang et al. [171] applied LoRA to fine-tune the Wav2vec2 model [8], achieving comparable performance to full fine-tuning while reducing trainable parameters by a factor of 198 compared with the full fine-tuning. For multilingual automatic speech recognition (ASR), Xu et al. [172] proposed O-LoRA and O-AdaLoRA to adapt Whisper model [170] for low-resource languages like Uyghur and Tibetan. Similarly, Song et al [173] proposed LoRA-Whisper, that incorporates LoRA into Whisper, effectively mitigating language interference and facilitating the addition of new languages without degrading existing performance. Liu et al. [174] introduce LoRA-Whisper which extends LoRA to low-resource ASR by comparing it with bottleneck adapters for fine-tuning Whisper across seven low-resource languages.
@@ 5.4 LoRA in Code Engineering
For code generation and summarization, Kumar et al. [177] developed LoRA-based federated learning methods for code summarization, which preserves data privacy without direct access to source code. Cui et al. [178] introduced OriGen, which leverages code-to-code augmentation and self-reflection techniques for generating high-quality Register Transfer Level (RTL)222RTL is a design abstraction that describes digital circuits in terms of data flow between registers and their logical operations. code.
@@ 5.5 LoRA in Scientific Discovery
LoRA has been utilized across a wide range of scientific fields, including molecular tasks [179, 180, 96, 181, 182], and materials science [96].
In the field of protein analysis, Zeng et al.[179] developed PEFT-SP, a framework utilizing LoRA to fine-tune large protein language models (PLMs) like ESM-2 [183] for signal peptide prediction. This approach significantly improved performance, especially for rare peptide types, while mitigating overfitting and maintaining low computational costs. Similarly, Schmirler et al.[180] applied LoRA to PLMs, like like ProtT5 [184] and ESM-2, across various protein-related tasks, demonstrating accelerated training and improved downstream predictions while preventing catastrophic forgetting. In another molecular application, Schreiber [181] employed LoRA and its quantized version, QLoRA [55], in the ESMBind and QBind models for protein binding site and post-translational modification predictions. These models achieved enhanced generalization on unseen protein sequences without relying on structural data or multiple sequence alignments. Lv et al.[185] introduced ProLLaMA, demonstrating that a two-stage LoRA approach with different ranks could enable effective protein language learning while maintaining natural language capabilities.
In more general science area, Buehler et al. [96] developed X-LoRA, a framework that dynamically combines multiple specialized LoRA adapters. This approach demonstrated exceptional capability in solving inverse and forward tasks in protein mechanics and material design, illustrating LoRA’s potential for complex multidisciplinary applications.
@@ 5.6 LoRA in Recommender Systems
For sequential recommendation, Qin et al. [188] introduced ATFLRec which integrates audio and text data by independently optimizing LoRA modules for each modality. Zheng et al. [189] developed LLM-TRSR for text-rich recommendation scenarios, leveraging LoRA’s efficient fine-tuning to handle large-scale text data while maintaining real-time capabilities. Kong et al. [190] developed iLoRA that employs an MoE framework, creating specialized LoRA modules for distinct user behavior patterns, dynamically engaging relevant experts based on individual interaction sequences. Recently, Ji et al.[191] proposed GenRec, a pure text-based LLM that uses item names as IDs and leverages LoRA to fine-tune LLaMA, showing superior performance on large-scale datasets where LLMs can directly learn collaborative information from text representations.
@@ 5.7 LoRA in Graph Learning
Recently, researchers have also explored the applications of LoRA on non-Euclidean data (i.e., graphs), fine-tuning graph neural networks (GNN) [192] to adapt to new graphs or structure updates of existing graphs.
For cross-domain graph neural network adaption, Yang et al.[193] introduced GraphLoRA, which constructs a small trainable GNN alongside the pre-trained one to bridge structural and feature distribution gaps between different graphs. By applying low-rank decomposition to the tunable GNN parameters and incorporating graph structure-based regularization, GraphLora can effectively adapt pre-trained GNNs to diverse graph domains by fine-tuning only 20% of the parameters.
@@ 5.8 LoRA in Spatial-Temporal Forecasting
Multivariate time series data is prevalent in real-world scenarios such as transportation, weather forecasting, and economics [195]. Recent studies have explored using LoRA to address specific challenges in this field, such as node-specific adaptation and multi-channel modeling and out-of-domain prediction.
@@
For out-of-domain prediction, Gupta et al.[198] analyzed the effectiveness of LoRA-based fine-tuning across leading time series foundation models like Lag-Llama[199], MOIRAI [200] and Chronos [201], showing improved vital sign of sepsis patients in intensive care units forecasting while reducing computational costs. Ren et al.[202] introduced TPLLM, which injects trainable rank-decomposition matrices into GPT-2 Transformer blocks for traffic forecasting, effectively adapting the model to process spatial-temporal representations with limited historical traffic data.
@@ 5.9 LoRA in Multi-Modal
Multimodal Foundation Models (MFMs) combine different data modalities including text, audio, images, video, etc within a shared representational space, enabling cross-modal reasoning and understanding. LoRA enhances these models by optimizing training efficiency while strengthening inter-modal alignments. Language-vision and language-audio learning represent two main domains where LoRA has been extensively applied to adapt MFMs.

### v2

@@ 4.5 LoRA for Long Sequence Modeling
The ability to process long sequences is crucial for many tasks across various domains handled by foundation models [144, 145, 146]. However, standard foundation models are typically constrained by their maximum context length due to the quadratic computational complexity of self-attention with respect to sequence length. To address this limitation, several LoRA-based techniques have been proposed to extend the context window of foundation models.
@@ 4.6 LoRA Serving Systems
S-LoRA [148] introduced a unified paging mechanism to manage both KV cache and LoRA weights in GPU memory, enabling concurrent serving of thousands of LoRA adapters. Punica [149] developed a custom CUDA kernel, Segmented Gather Matrix-Vector Multiplication (SGMV), facilitating efficient batching of requests across different LoRA models on a single GPU. CARASERVE [150] adopted a CPU-assisted strategy, initiating prefill computations for newly requested adapters during GPU loading to mitigate cold-start latency. Furthermore, CARASERVE introduced a rank-aware scheduling algorithm for optimized request routing in multi-GPU clusters. CA-LoRA [151] incorporated LoRA knowledge inheritance and model knowledge recovery mechanisms to maintain performance on personal devices. These innovations enhance the scalability and efficiency of serving multiple LoRA adapters across diverse computing environments.
5 Applications
The effectiveness and efficiency of LoRA in fine-tuning foundation models have led to its widespread adoption across domains, including language processing, computer vision, speech recognition, multimodal learning, code engineering, scientific discovery, recommender systems, graph learning, and spatiotemporal forecasting.
@@ 5.1 LoRA in Language Tasks
Language foundation models such as LLaMA [4], RoBERTa [21], and DeBERTa [152] serve as essential bases in LoRA research and have been extensively studied across tasks including natural language understanding [48, 46, 49], question answering [36, 153], machine translation [50, 46], reasoning [56, 50, 82, 154], and natural language generation [55, 64, 65]. This section highlights LoRA’s applications in specialized NLP domains.
Multilingual Language and Dialect Processing. LoRA enables efficient multilingual adaptation while preserving base models’ capabilities through minimal parameter updates. FedLFC [155] introduces a LoRA-based fine-tuning framework that uses pretrained language-specific LoRA experts, combining them into a mixture of language experts (MoLE) and enhancing performance through knowledge distillation for efficient speech recognition. LAMPAT [156] leverages LoRA for unsupervised multilingual paraphrasing by applying adversarial training to generate diverse outputs while preserving semantic meaning across languages. HyperLoRA [157] generates dialect-specific LoRA adapters using linguistic feature vectors, enabling zero-shot adaptation to unseen English dialects without requiring dialectal training data.
Medical and Clinical Text Processing. Medical and clinical text processing faces unique challenges due to limited data availability and the sensitive nature of clinical datasets. LoRA has emerged as a promising solution to address these limitations. Ji et al. [158] improved assertion detection with minimal data in clinical settings, while Le et al. [159] demonstrated LoRA’s adaptability for clinical NLP tasks in data-limited environments. Liu et al. [104] combined LoRA with MoE in a multi-task framework to tackle data imbalance. Shi et al. [160] used LoRA in the MedAdapter framework for test-time adaptation, minimizing computational costs and data sharing. Christophe et al. [161] proposed the Med42 model, showing that LoRA fine-tuning outperforms traditional methods on key medical benchmarks such as the USMLE.
@@ 5.2 LoRA in Computer Vision
LoRA has been effectively applied to vision foundation models such as ViTs [165], DINOv2 [166], MAE [6], SAM [7], and Florence [167], enhancing adaptability across visual understanding and visual generation tasks.
@@ 5.2.1 Visual Understanding
Visual understanding contains a broad spectrum of tasks, including domain adaptation, semantic segmentation, and content authenticity checking.
@@
ExPLoRA [168] integrates LoRA modules into ViTs’ self-attention mechanisms, capturing domain-specific style variations, such as in satellite imagery. Similarly, MeLo [169] applies LoRA for fine-tuning in medical imaging, such as thoracic disease classification. Additionally, Kong et al. [170] use LoRA to improve the generalization of ViTs for face forgery detection across various manipulation techniques and datasets.
A notable advancement in this direction is ConvLoRA [171], which extends the LoRA paradigm to convolutional neural networks for unsupervised domain adaptation in medical images. The architecture combines trainable low-rank decomposition matrices with adaptive batch normalization, establishing a more robust framework for domain transfer.
Semantic Segmentation. The adaptation of visual foundation models, particularly the Segment Anything Model (SAM) [7], has witnessed significant progress in semantic segmentation through LoRA integration. ConvLoRA [171] enhances SAM [7] for semantic segmentation in remote sensing, medical, and agricultural images. Building upon this foundation, SAMed [172] demonstrates the effectiveness of LoRA-based fine-tuning for multi-organ segmentation tasks. SurgicalSAM [173] applies similar techniques to the domain of robotic surgical instrument segmentation.
Content Authenticity Checking. Detecting synthesized content has become increasingly important with the advancement of generative models. CLIPMoLE [174] combines shared and separate LoRA modules within an MoE framework to efficiently fine-tune CLIP ViT models for transferable image detection across diverse generation techniques. Similarly, MoE-FFD [175] integrates LoRA with convolutional adapter modules, targeting face forgery detection while keeping the ViT backbone frozen.
@@ 5.2.2 Visual Generation
Image Stylization. LoRA has become a key technique for image stylization, enabling rapid adaptation of diffusion models to specific artistic styles while preserving the base model’s generative diversity. [179] employ LoRA to efficiently fine-tune Stable Diffusion [10] for comic-style transfer, specifically targeting the style of Calvin and Hobbes comics. Frenkel et al. [180] propose B-LoRA, which leverages Stable Diffusion XL [181] to implicitly disentangle style and content from a single image, enabling tasks such as style transfer and text-guided stylization. Borse et al. [182] introduce FouRA, a frequency-domain low-rank adaptation method that mitigates distribution collapse and data copying issues common in standard LoRA fine-tuning. Recently, DragLoRA [183] integrates LoRA adapters into drag-based editing to enhance precision and efficiency in foreground object manipulation.
Multi-Concept Customization. LoRA is widely adopted for multi-concept customization in visual generation, where models must synthesize images combining multiple subjects or styles. Shah et al. [184] propose ZipLoRA, which merges independently trained LoRAs via embedding-decomposed adaptation and gradient fusion while preserving concept identities. Gu et al. [185] and Zhong et al [186] explore LoRA switching and composition to enhance diffusion control and multi-element generation, respectively. Yang et al. [187] introduce LoRA-Composer, a training-free framework that integrates multiple LoRAs through concept injection and isolation to mitigate concept vanishing and confusion. Po et al. [188] enforce orthogonal parameter updates to enable efficient merging of concept-specific LoRAs. MoLE [189] employs low-rank experts with customized datasets for human-centric generation, while Zhuang et al. [190] propose the TimeStep Master framework, assigning distinct LoRA experts to different diffusion timesteps.
@@
Whether addressing multi-concept synthesis, high-resolution tasks, style transfer, or continual learning, LoRA provides a lightweight solution enabling models like Stable Diffusion and CLIP to maintain high performance while being adapted for new and complex tasks.
@@ 5.3 LoRA in Speech Recognition
LoRA has been widely used in speech recognition tasks, particularly in efficiently fine-tuning foundation models like Wav2vec2 [8] and Whisper [193].
In fake audio detection, Wang et al. [194] applied LoRA to fine-tune Wav2vec2 [8], achieving performance comparable to full fine-tuning while reducing trainable parameters by a factor of 198. For multilingual automatic speech recognition (ASR), Xu et al. [195] proposed O-LoRA and O-AdaLoRA to adapt Whisper [193] for low-resource languages such as Uyghur and Tibetan. Similarly, Song et al. [196] introduced LoRA-Whisper, which integrates LoRA into Whisper to mitigate language interference and add new languages without degrading existing performance. Liu et al. [197] further extended LoRA-Whisper to low-resource ASR, demonstrating advantages over bottleneck adapters when fine-tuning Whisper across seven low-resource languages.
@@ 5.4 LoRA in Code Engineering
For code generation and summarization, Kumar et al. [200] developed LoRA-based federated learning methods for code summarization, preserving data privacy without direct access to source code. Cui et al. [201] introduced OriGen, which leverages code-to-code augmentation and self-reflection techniques for generating high-quality register-transfer-level (RTL)222RTL is a design abstraction that describes digital circuits in terms of data flow between registers and their logical operations. code.
@@ 5.5 LoRA in Scientific Discovery
LoRA has been utilized across a wide range of scientific fields, including molecular tasks and materials science.
In protein analysis, Zeng et al. [202] developed PEFT-SP, a framework that leverages LoRA to fine-tune large protein language models (PLMs) such as ESM-2 [203] for signal peptide prediction. This approach significantly improved performance, especially for rare peptide types, while mitigating overfitting and maintaining low computational costs. Similarly, Schmirler et al. [204] applied LoRA to PLMs such as ProtT5 [205] and ESM-2 across diverse protein-related tasks, demonstrating accelerated training and improved downstream predictions while preventing catastrophic forgetting. In another molecular application, Schreiber [206] employed LoRA and its quantized variant, QLoRA [55], in the ESMBind and QBind models for protein binding-site and post-translational modification prediction. These models achieved enhanced generalization on unseen protein sequences without relying on structural data or multiple sequence alignments. Lv et al. [207] introduced ProLLaMA, demonstrating that a two-stage LoRA approach with varying ranks enables effective protein language learning while maintaining natural language capabilities.
Beyond molecular science, Buehler et al. [108] developed X-LoRA, a framework that dynamically combines specialized LoRA adapters to solve inverse and forward problems in protein mechanics and material design. Dagdelen et al. [208] utilized LoRA to enhance scientific information extraction, enabling more effective data processing and knowledge discovery. These methods highlight LoRA’s potential for complex multidisciplinary applications.
@@ 5.6 LoRA in Recommender Systems
For sequential recommendation, Qin et al. [211] introduced ATFLRec, which integrates audio and text data by independently optimizing LoRA modules for each modality. Zheng et al. [212] developed LLM-TRSR for text-rich recommendation scenarios, leveraging LoRA’s efficient fine-tuning to handle large-scale text data while maintaining real-time responsiveness. Kong et al. [213] proposed iLoRA, which employs an MoE framework with specialized LoRA modules for distinct user behavior patterns, dynamically engaging relevant experts based on individual interaction sequences. Recently, Ji et al. [214] introduced GenRec, a text-centric LLM that treats item names as identifiers and fine-tunes LLaMA with LoRA, achieving superior performance on large-scale datasets by directly learning collaborative information from textual representations.
@@ 5.7 LoRA in Graph Learning
Recent work has also explored applying LoRA to non-Euclidean data (i.e., graphs) by fine-tuning graph neural networks (GNNs) [215] to adapt to new graphs or structural updates of existing ones.
For cross-domain graph neural network adaptation, Yang et al. [216] introduced GraphLoRA, which constructs a small trainable GNN alongside the pretrained model to bridge structural and feature distribution gaps between graphs. By applying low-rank decomposition to the tunable GNN parameters and incorporating graph-structure-based regularization, GraphLoRA adapts pretrained GNNs to diverse domains while fine-tuning only 20% of the parameters.
@@
For equivariant GNNs, ELoRA [218] is tailored to SO(3)-equivariant architectures, maintaining physical symmetry during fine-tuning and improving the accuracy of molecular energy and force predictions.
@@ 5.8 LoRA in Spatial-Temporal Forecasting
Multivariate time series data is prevalent in real-world scenarios such as transportation, weather forecasting, and economics [219, 220, 221]. Recent studies have explored using LoRA to address challenges in this field, including node-specific adaptation, multi-channel modeling, and out-of-domain prediction.
@@
For out-of-domain prediction, Gupta et al. [224] analyzed the effectiveness of LoRA-based fine-tuning across leading time-series foundation models such as Lag-Llama [225], MOIRAI [226], and Chronos [227], demonstrating improved forecasting of sepsis patients’ vital signs in intensive care units while reducing computational costs. Ren et al.[228] introduced TPLLM, which injects trainable rank-decomposition matrices into GPT-2 Transformer blocks for traffic forecasting, effectively adapting the model to process spatial-temporal representations with limited historical traffic data.
@@ 5.9 LoRA in Multi-Modal
Multimodal foundation models (MFMs) combine modalities such as text, audio, images, and video within a shared representational space, enabling cross-modal reasoning and understanding. LoRA enhances these models by improving training efficiency while strengthening inter-modal alignment. Language-vision and language-audio learning represent two primary domains where LoRA has been extensively applied to adapt MFMs.